# Accurate Power Spectrum Estimation toward Nyquist limit

### v1

@@ 4 Conclusion
In this article, we have presented four approaches for power spectrum calculation using FFT and compared their performance. It is widely known by the community that the shot noise is a constant $\frac{1}{N}$ as halo/galaxy distribution is Poissonian. Although this approximation is widely adopted in general estimations (as Eq. (2.10) and Eq. (2.17)), when using the interlacing technique, the aliasing effects on the signal term are reduced, and thus the impact of aliasing on the shot noise term can no longer be neglected. We have provided the specific form of the shot noise term under the interlacing technique in Eq. (2.15), which demonstrates how the shot noise term is affected by the window function and aliasing effects. We have also pointed out that Eq. (2.16) provides a better estimation compared to Eq. (2.17). In precise cosmological analysis, this algorithm allows for a more accurate subtraction of shot noise, helping to reduce the systematic bias induced by the detailed numerical calculation. This reduction of the systematics toward the Nyquist frequency also saves the computational cost, since for a good power spectrum estimator such as Eq. (2.16) we could use coarse mesh in the analysis for a given precision requirement on a specified scale. This is of great importance as the precise cosmological constraints rely on large suite of mock observation of order $\mathcal{O}(10^{4})$ to obtain the reasonably good covariance matrix [12]. Moreover, as realized in literature, the higher-order window functions yield better accuracy, particularly under the interlacing technique. Considering the increased computational cost associated with higher-order window functions, it is not recommended to increase the order of the window function without adopting the interlacing technique, as it does not effectively reduce systematic errors. Under similar consideration, using better estimator could save computational cost by only using low order window function.
@@ Acknowledgments
This work is supported by the National Key R&D Program of China (2023YFA1607800, 2023YFA1607802), the National Science Foundation of China (grant numbers 12273020), the China Manned Space Project with numbers CMS-CSST-2021-A03, the ‘111’ Project of the Ministry of Education under grant number B20019, and the sponsorship from Yangyang Development Fund. This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
@@ Appendix A Isotropic Approximation for Shot Noise
Figure 7: Analytical shot noise expressions are presented. On the left side are the results for the CIC window function, and on the right side are the results for the TSC window function. The dashed lines are the isotropic version following Eq. (A.2) for interlacing and Eq. (2.8) for non-interlacing. The solid lines are the exact calculation, following Eq. (2.15) and Eq. (A.1) for interlacing and non-interlacing, respectively. The scattered points are the first term in Eq. (2.15).
@@
$C^{\mathrm{iso}}(k)=\left\{\begin{aligned} &1\ ,&\text{NGP}\ ,\\ &1-\frac{2}{3}\sin^{2}\left(\frac{\pi k}{k_{H}}\right)\ ,&\text{CIC}\ ,\\ &1-\sin^{2}\left(\frac{\pi k}{k_{H}}\right)+\frac{2}{15}\sin^{4}\left(\frac{\pi k}{k_{H}}\right)\ ,&\text{TSC}\ ,\\ &1-\frac{4}{3}\sin^{2}\left(\frac{\pi k}{k_{H}}\right)+\frac{2}{5}\sin^{4}\left(\frac{\pi k}{k_{H}}\right)-\frac{4}{315}\sin^{6}\left(\frac{\pi k}{k_{H}}\right)\ ,&\text{PSC}\ .\\ \end{aligned}\right.$ (A.1)
In short, it approximates $\prod_{i}S_{2p}[\sin^{2}(\pi k_{i}/k_{H})]$ as $S_{2p}[\sin^{2}(\pi k/k_{H})]$. For the case with interlacing, the exact anisotropic shot noise is Eq. (2.15). The similar approximation above can only be straightforwardly applied to the first term $\prod_{i}C_{e}(k_{i})$. We show that the first term is overwhelming in Figure 7. Keeping the first term only gives an relative error about 0.3% when approaching Nyquist frequency with CIC window function, and 0.02% with TSC window function.
Given that the above approximation is valid, it is straightforward to treat the isotropic version of the first term $\prod_{i}C_{e}(\pi k_{i}/2k_{H})\approx C_{e}(\pi k/2k_{H})$ as the isotropic approximation for interlacing.
$C^{{\operatorname{int}},\mathrm{iso}}(k)=\left\{\begin{aligned} &\cos^{2}\left(\frac{\pi k}{2k_{H}}\right),&\text{NGP}\ ,\\ &\cos^{4}\left(\frac{\pi k}{2k_{H}}\right)\left[1-\frac{2}{3}\sin^{2}\left(\frac{\pi k}{2k_{H}}\right)\right]\ ,&\text{CIC}\ ,\\ &\cos^{6}\left(\frac{\pi k}{2k_{H}}\right)\left[1-\sin^{2}\left(\frac{\pi k}{2k_{H}}\right)+\frac{2}{15}\sin^{4}\left(\frac{\pi k}{2k_{H}}\right)\right]\ ,&\text{TSC}\ ,\\ &\cos^{8}\left(\frac{\pi k}{2k_{H}}\right)\left[1-\frac{4}{3}\sin^{2}\left(\frac{\pi k}{2k_{H}}\right)+\frac{2}{5}\sin^{4}\left(\frac{\pi k}{2k_{H}}\right)-\frac{4}{315}\sin^{6}\left(\frac{\pi k}{2k_{H}}\right)\right]\ ,&\text{PSC}\ .\\ \end{aligned}\right.$ (A.2)
In short, $C^{{\operatorname{int}},{\mathrm{iso}}}(k)=\cos^{2p}(\pi k/2k_{H})C^{\mathrm{iso}}(k/2)$. Note that the prefactor of $\cos^{2p}x$ can be expressed as polynomial of $\sin^{2}x$ up to order $p$, leaving the whole formula only containing sine terms. Or we can express all the terms in the polynomial of $\sin^{2}x$ into polynomial of $\cos^{2}x$. However, these further expansions help less when we input the above formula using programming language. Thus we keep this mixed form of cosine and sine terms in $C^{{\operatorname{int}},{\mathrm{iso}}}(k)$. Also in Figure 7, the exact shot noise term is compared with the isotropic shot noise for both interlacing and non-interlacing case. The isotropic approximation for interlacing case is better than the one without interlacing.
@@ B.1 Equal-spacing interlacing scheme
$=\sum_{\boldsymbol{n}}{\theta_{\boldsymbol{n}}{\delta}^{f}({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})}\ .$
@@
$C^{\operatorname{int-1D}}_{m}(k)$ $=|W(k)|^{2}\sum_{n=ml}\left(\frac{k}{k-nk_{H}}\right)^{2p}=|W(k)|^{2}\sum_{l}\left(\frac{k}{k-mlk_{H}}\right)^{2p}$ (B.4)
$=|W(k)|^{2}\sum_{l}\left(\frac{k/m}{k/m-lk_{H}}\right)^{2p}=\frac{|W(k)|^{2}}{|W(k/m)|^{2}}C(k/m)$
$=\frac{\sin^{2p}\left(\frac{\pi k}{k_{H}}\right)}{m^{2p}\sin^{2p}\left(\frac{\pi k}{mk_{H}}\right)}S_{2p}\left[\sin^{2}\left(\frac{\pi k}{mk_{H}}\right)\right]\ .$
@@
$C^{\operatorname{int-1D},n_{\textrm{shift}}}_{m}(k)$ $=|W(k)|^{2}\sum_{l}\left(\frac{k}{k-(ml+n_{\textrm{shift}})k_{H}}\right)^{2p}$ (B.5)
$=\frac{\sin^{2p}\left(\frac{\pi k}{k_{H}}\right)}{m^{2p}\sin^{2p}\left(\frac{\pi k}{mk_{H}}-\frac{\pi n_{\textrm{shift}}}{m}\right)}S_{2p}\left[\sin^{2}\left(\frac{\pi k}{mk_{H}}-\frac{\pi n_{\textrm{shift}}}{m}\right)\right]\ .$
@@ B.2 Bisection interlacing scheme
We list the $\theta_{\boldsymbol{n}}$ for using $m=1$ (i.e., no interlacing), $2$, $4$, $8$ and $16$ interlaced grids.
@@
The derivation of shot noise for $m=2^{3h}$ interlaced grids is straightforward because the summation of each direction can be done individually.
$C^{\operatorname{int}}_{2^{3h}}({\boldsymbol{k}})$ $=W^{2}({\boldsymbol{k}})\frac{\sum_{{\boldsymbol{n}}}\theta_{\boldsymbol{n}}|W({\boldsymbol{k}}+k_{H}{\boldsymbol{n}})|^{2}}{W^{2}({\boldsymbol{k}})}$ (B.9)
$=W^{2}({\boldsymbol{k}})\prod_{i=x,y,z}\sum_{{\boldsymbol{n}}}{\prod_{j=1}^{h}}\frac{1}{2}\left[1+{e^{-\frac{\pi\mathrm{i}n_{i}}{2^{j-1}}}}\right]\left(\frac{k_{i}}{k_{i}-n_{i}k_{H}}\right)^{2p}$
@@
$C^{\operatorname{int-1D,even}}_{2^{h}}(k)=C^{\operatorname{int-1D}}_{2^{h-1}}(k)\ .$ (B.10)
@@
$C^{\operatorname{int-1D,odd}}_{m=2^{h}}(k)$ $=|W(k)|^{2}\sum_{n}\left(\frac{k}{k-(2^{h}n+2^{h-1})k_{H}}\right)^{2p}$ (B.11)
$=\frac{\sin^{2p}\left(\frac{\pi k}{k_{H}}\right)}{m^{2p}\cos^{2p}\left(\frac{\pi k}{mk_{H}}\right)}S_{2p}\left[\cos^{2}\left(\frac{\pi k}{mk_{H}}\right)\right]\ .$
@@
$C^{\operatorname{int-1D,odd}}_{2^{h}}(k)=C^{\operatorname{int-1D}}_{2^{h}}(k)-C^{\operatorname{int-1D,even}}_{2^{h}}(k)=C^{\operatorname{int-1D}}_{2^{h}}(k)-C^{\operatorname{int-1D}}_{2^{h-1}}(k)\ .$ (B.12)
@@
The last point worth mentioning is that $C_{8}^{\operatorname{int}}({\boldsymbol{k}})/W^{2}({\boldsymbol{k}})$ for 8 interlaced grids is exactly same as $C({\boldsymbol{k}})/W^{2}({\boldsymbol{k}})$ with no interlacing but doubling the resolution. In principle, at least for NGP, given the density fields on 8 interlaced grids, the density on two times finer grid can be solved out. This explains the above equivalence. The computational cost is roughly same, but the interlacing technique allows lower memory usage.

### v2

@@ 4 Conclusion
In this article, we have presented four approaches for power spectrum calculation using FFT and compared their performance. It is widely known by the community that the shot noise is a constant $\frac{1}{N}$ as halo/galaxy distribution is Poissonian. Although this approximation is widely adopted in general estimations (as Eq. (2.10) and Eq. (2.17)), when using the interlacing technique, the aliasing effects on the signal term are reduced, and thus the impact of aliasing on the shot noise term can no longer be neglected. We have provided the specific form of the shot noise term under the interlacing technique in Eq. (2.15), which demonstrates how the shot noise term is affected by the window function and aliasing effects.
In precise cosmological analysis, the above accurate shot noise form allows for a more accurate noise subtraction from the measurement, helping to reduce the systematic bias induced by the detailed numerical calculation. We have pointed out that Eq. (2.16) provides a better estimation compared to Eq. (2.17). For the three dark matter samples with number density ranging from $10^{-2}\text{ Mpc}^{-3}h^{3}$ to $10^{-4}\text{ Mpc}^{-3}h^{3}$ analyzed on a coarse mesh of size $3\,h^{-1}{\rm Mpc}$ with CIC window, the systematic bias from Eq. (2.16) is well under control up to the Nyquist frequency. On the contrary, the bias induced by the estimator Eq. (2.17) that ignore the aliasing on the shot noise typically exceeds the statistical uncertainty on the frequency beyond 0.85 times the Nyquist frequency. For the cases adopting TSC window, the induced bias from the two estimators is well below the statistical uncertainty. But Eq. (2.16) always shows lower bias than Eq. (2.17).
This reduction of the systematics toward the Nyquist frequency also saves the computational cost, since for a good power spectrum estimator such as Eq. (2.16) we could use coarse mesh in the analysis for a given precision requirement on a specified scale. This is of great importance as the precise cosmological constraints rely on large suite of mock observation of order $\mathcal{O}(10^{4})$ to obtain the reasonably good covariance matrix [12]. Moreover, as realized in literature, the higher-order window functions yield better accuracy, particularly under the interlacing technique. Considering the increased computational cost associated with higher-order window functions, it is not recommended to increase the order of the window function without adopting the interlacing technique, as it does not effectively reduce systematic errors. Under similar consideration, using better estimator could save computational cost by only using low order window function.
@@ Acknowledgments
We thank Zhao Chen for the useful discussions and the anonymous referee for the valuable suggestions. This work is supported by the National Key R&D Program of China (2023YFA1607800, 2023YFA1607802), the National Science Foundation of China (grant numbers 12273020), the China Manned Space Project with numbers CMS-CSST-2021-A03, the ‘111’ Project of the Ministry of Education under grant number B20019, and the sponsorship from Yangyang Development Fund. This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
@@ Appendix A Isotropic approximation for shot noise
Figure 8: Analytical shot noise expressions are presented. On the left side are the results for the CIC window function, and on the right side are the results for the TSC window function. The dashed lines are the isotropic version following Eq. (A.2) for interlacing and Eq. (2.8) for non-interlacing. The solid lines are the exact calculation, following Eq. (2.15) and Eq. (A.1) for interlacing and non-interlacing, respectively. The data points are the first term in Eq. (2.15).
@@
$C^{\mathrm{iso}}(k)=\left\{\begin{aligned} &1\ ,&\text{NGP}\ ,\\ &1-\frac{2}{3}\sin^{2}\left(\frac{\pi k}{2k_{N}}\right)\ ,&\text{CIC}\ ,\\ &1-\sin^{2}\left(\frac{\pi k}{2k_{N}}\right)+\frac{2}{15}\sin^{4}\left(\frac{\pi k}{2k_{N}}\right)\ ,&\text{TSC}\ ,\\ &1-\frac{4}{3}\sin^{2}\left(\frac{\pi k}{2k_{N}}\right)+\frac{2}{5}\sin^{4}\left(\frac{\pi k}{2k_{N}}\right)-\frac{4}{315}\sin^{6}\left(\frac{\pi k}{2k_{N}}\right)\ ,&\text{PSC}\ .\\ \end{aligned}\right.$ (A.1)
In short, it approximates $\prod_{i}S_{2p}[\sin^{2}(\pi k_{i}/2k_{N})]$ as $S_{2p}[\sin^{2}(\pi k/2k_{N})]$. For the case with interlacing, the exact anisotropic shot noise is Eq. (2.15). A similar approximation can only be straightforwardly applied to the first term $\prod_{i}C_{e}(k_{i})$. We show that the first term is overwhelming in Figure 8. Keeping the first term only gives an relative error about 0.3% when approaching Nyquist frequency with CIC window function, and 0.02% with TSC window function.
Given that the above approximation is valid, it is straightforward to treat the isotropic version of the first term $\prod_{i}C_{e}(\pi k_{i}/4k_{N})\approx C_{e}(\pi k/4k_{N})$ as the isotropic approximation for interlacing,
$C^{{\operatorname{int}},\mathrm{iso}}(k)=\left\{\begin{aligned} &\cos^{2}\left(\frac{\pi k}{4k_{N}}\right),&\text{NGP}\ ,\\ &\cos^{4}\left(\frac{\pi k}{4k_{N}}\right)\left[1-\frac{2}{3}\sin^{2}\left(\frac{\pi k}{4k_{N}}\right)\right]\ ,&\text{CIC}\ ,\\ &\cos^{6}\left(\frac{\pi k}{4k_{N}}\right)\left[1-\sin^{2}\left(\frac{\pi k}{4k_{N}}\right)+\frac{2}{15}\sin^{4}\left(\frac{\pi k}{4k_{N}}\right)\right]\ ,&\text{TSC}\ ,\\ &\cos^{8}\left(\frac{\pi k}{4k_{N}}\right)\left[1-\frac{4}{3}\sin^{2}\left(\frac{\pi k}{4k_{N}}\right)+\frac{2}{5}\sin^{4}\left(\frac{\pi k}{4k_{N}}\right)-\frac{4}{315}\sin^{6}\left(\frac{\pi k}{4k_{N}}\right)\right]\ ,&\text{PSC}\ .\\ \end{aligned}\right.$ (A.2)
In short, $C^{{\operatorname{int}},{\mathrm{iso}}}(k)=\cos^{2p}(\pi k/4k_{N})C^{\mathrm{iso}}(k/2)$. Note that the prefactor of $\cos^{2p}x$ can be expressed as polynomial of $\sin^{2}x$ up to order $p$, leaving the whole formula only containing sine terms. Or we can express all the terms in the polynomial of $\sin^{2}x$ into polynomial of $\cos^{2}x$. However, these further expansions help less when we input the above formula using programming language. Thus we keep this mixed form of cosine and sine terms in $C^{{\operatorname{int}},{\mathrm{iso}}}(k)$. Also in Figure 8, the exact shot noise term is compared with the isotropic shot noise for both interlacing and non-interlacing case. The isotropic approximation for interlacing case is better than the one without interlacing.
@@ Appendix B Shot noise for high order interlacing
Refer to caption Refer to caption
Figure 11: Shot noise estimated from Poisson-distributed random samples with high order interlacing under CIC window function. On the left side are the results for the equal-spacing interlacing scheme with 2, 3 and 4 interlaced grids, and on the right side are the results for the bisection interlacing scheme with 2, 4 and 8 interlaced grids. For the top panels the data points represent the $\langle\delta^{f}(k)^{2}\rangle/W^{2}(k)$ obtained from the random sample, the dot-dashed lines and the solid lines represent the theoretical formulae given by Eq. (B.6), Eq. (B.9), Eq. (B.13) and Eq. (B.14), respectively. The bottom panels show the deviation between the individual results and their respective theoretical formula. The analytical expressions describe the shot noise accurately, leaving only the same sample variance scattering around unity.
@@ B.1 Equal-spacing interlacing scheme
$=\sum_{\boldsymbol{n}}{\theta_{\boldsymbol{n}}{\delta}^{f}({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})}\ .$
@@
$C^{\operatorname{int-1D}}_{m}(k)$ $=|W(k)|^{2}\sum_{n=ml}\left(\frac{k}{k-2nk_{N}}\right)^{2p}=|W(k)|^{2}\sum_{l}\left(\frac{k}{k-2mlk_{N}}\right)^{2p}$ (B.4)
$=|W(k)|^{2}\sum_{l}\left(\frac{k/m}{k/m-2lk_{N}}\right)^{2p}=\frac{|W(k)|^{2}}{|W(k/m)|^{2}}C(k/m)$
$=\frac{\sin^{2p}\left(\frac{\pi k}{2k_{N}}\right)}{m^{2p}\sin^{2p}\left(\frac{\pi k}{2mk_{N}}\right)}S_{2p}\left[\sin^{2}\left(\frac{\pi k}{2mk_{N}}\right)\right]\ .$
@@
$C^{\operatorname{int-1D},n_{\textrm{shift}}}_{m}(k)$ $=|W(k)|^{2}\sum_{l}\left(\frac{k}{k-2(ml+n_{\textrm{shift}})k_{N}}\right)^{2p}$ (B.5)
$=\frac{\sin^{2p}\left(\frac{\pi k}{2k_{N}}\right)}{m^{2p}\sin^{2p}\left(\frac{\pi k}{2mk_{N}}-\frac{\pi n_{\textrm{shift}}}{m}\right)}S_{2p}\left[\sin^{2}\left(\frac{\pi k}{2mk_{N}}-\frac{\pi n_{\textrm{shift}}}{m}\right)\right]\ .$
@@ B.2 Bisection interlacing scheme
We list the $\theta_{\boldsymbol{n}}$ for using $m=1$ (i.e., no interlacing), $2$, $4$, $8$ and $16$ interlaced grids,
@@
The derivation of shot noise for $m=2^{3h}$ interlaced grids is straightforward because the summation of each direction can be done individually,
$C^{\operatorname{int}}_{2^{3h}}({\boldsymbol{k}})$ $=W^{2}({\boldsymbol{k}})\frac{\sum_{{\boldsymbol{n}}}\theta_{\boldsymbol{n}}|W({\boldsymbol{k}}+2k_{N}{\boldsymbol{n}})|^{2}}{W^{2}({\boldsymbol{k}})}$ (B.9)
$=W^{2}({\boldsymbol{k}})\prod_{i=x,y,z}\sum_{{\boldsymbol{n}}}{\prod_{j=1}^{h}}\frac{1}{2}\left[1+{e^{-\frac{\pi\mathrm{i}n_{i}}{2^{j-1}}}}\right]\left(\frac{k_{i}}{k_{i}-2n_{i}k_{N}}\right)^{2p}$
@@
$C^{\operatorname{int-1D,even}}_{2^{h}}(k)=C^{\operatorname{int-1D}}_{2^{h+1}}(k)\ .$ (B.10)
@@
$C^{\operatorname{int-1D,odd}}_{m=2^{h}}(k)$ $=|W(k)|^{2}\sum_{n}\left(\frac{k}{k-(2^{h+2}n+2^{h+1})k_{N}}\right)^{2p}$ (B.11)
$=\frac{\sin^{2p}\left(\frac{\pi k}{2k_{N}}\right)}{{(2m)}^{2p}\cos^{2p}\left(\frac{\pi k}{4mk_{N}}\right)}S_{2p}\left[\cos^{2}\left(\frac{\pi k}{4mk_{N}}\right)\right]\ .$
@@
$C^{\operatorname{int-1D,odd}}_{2^{h}}(k)=C^{\operatorname{int-1D}}_{2^{h}}(k)-C^{\operatorname{int-1D,even}}_{2^{h}}(k)=C^{\operatorname{int-1D}}_{2^{h}}(k)-C^{\operatorname{int-1D}}_{2^{h+1}}(k)\ .$ (B.12)
@@
The last point worth mentioning is that $C_{8}^{\operatorname{int}}({\boldsymbol{k}})/W^{2}({\boldsymbol{k}})$ for 8 interlaced grids given by Eq. (B.9) is exactly same as $C({\boldsymbol{k}})/W^{2}({\boldsymbol{k}})$ with no interlacing but doubling the resolution. In principle, at least for NGP, given the density fields on 8 interlaced grids, the density on two times finer grid can be solved out. This explains the above equivalence. The computational cost is roughly same, but the interlacing technique allows lower memory usage.
Finally, we present the performance of the Eq. (B.6), Eq. (B.9), Eq. (B.13) and Eq. (B.14) in Figure 11, specifically for $m=2,3,4$ with the equal-spacing interlacing scheme, and $m=2,4,8$ with the bisection interlacing scheme. This numerical validation confirms that the above complicate theoretical derivation is correct.